How do you solve this question where can i find more like this one
-1 ≤ x+1 ≤ 3
-2
The -2 is under the x+1

Answer by Alan3354(69443)

-1 <= (x+1)/(-2) <= 3
Multiply by 2
-2 <= -(x+1) <= 6
Multiply by -1, that inverts the <
2 >= x+1 >= -6
Subtract 1
1 >= x >= -7
Put in ascending order
-7 <= x <= 1
----------------
You can look at solved problems on the site, or google inequalities.